User Tools

Site Tools


handover:c1706_v1706

c1706/v1706

Hobart 26m:

Disk VSN: HOB+0094

Date volume at beginning: 7221.148 GB

  • 00:00 UT - experiment started OK (Patrick)
  • 00:00 UT - Lots of windstows, check with 'grep wind c1706ho.log'
  • 02:55 UT - Telescope does not reach stow sometimes when windstowing. Keep an eye on the live page, and issue an extra 'source=stow' command if the elevation is not changing. (Patrick)
  • received multiple warning of windstows, error m5 -900:already off, and error f1 -1 previous source in this schedule not reached before new source was commanded (04:46:10 UT till 07:01:38 UT) (mas)
  • At 07:00:00 UT, Warren called and halted the schedule for Hobart 26m (mas)
  • Started recording again at 07:24:37 UT (mas)
  • Received Alarm:error f1 -1 previous source in this schedule not reached before new source was commanded at 10:31:30 UT,10:34:24 UT, 10:38:23 UT, 10:40:52 UT, 10:44:02 UT, 10:45:48 UT, 10:50:12 UT, 10:52:24 UT, 10:54:20 UT, 10:57:12 UT, 10:59:20 UT,11:02:06 UT, 11:09:44 UT, 11:18: 13 UT and 11:20:06 UT. The left and right drives are OFF. Hydraulics fault/OFF. Have tried remote reset via the v-desk terminal on newsmerd, but activity idle. Have informed Warren and he go to Mt Pleasant. (mas)
  • Halt the schedule at 11:21:44 UT (mas)
  • Missed scans from scan 337-1033 to 337-1146 (mas)
  • Start recording again at 11:50:09 UT (mas)
  • 1806 UT missed scan 337-1806b due to slow slew speed. (Bryn)
  • 1909 UT missed scan 337-1909 due to slow slew speed. (Bryn)
  • 2330 UT missed scan 337-2327 missed, probably same slow slewing (Liz)

Katherine 12m:

Disk VSN: USN+0224

Date volume at beginning: 0 GB

  • 00:00 UT - experiment started OK (Patrick)
  • 2334 UT - at 22:00 UT approx. delay difference appears to start drifting rapidly (linear-like drift), may have stabilised at ~ -0.335 in the last few scans. Will monitor until end of experiment (in about 20mins), then reconfigure dbbc in break. (Liz)

Yarragadee 12m:

Disk VSN: HOB+0074

Date volume at beginning: 4189.252 GB

  • 00:00:00 UT - experiment started OK (Patrick)
  • 1800 UT - Humidity Sensor giving nonsense readings again (Bryn)
/home/www/auscope/opswiki/data/pages/handover/c1706_v1706.txt · Last modified: 2017/12/03 23:35 by Bryn Emptage